# Bell Nonlocality from Wigner Negativity in Qudit Systems

### v2

@@ Abstract
Nonlocality is an essential concept that distinguishes quantum from classical models and has been extensively studied in systems of qubits. For higher-dimensional systems, certain results for their two-level counterpart, like Bell violations with stabilizer states and Clifford operators, do not generalize. On the other hand, similar to continuous variable systems, Wigner negativity is necessary for nonlocality in qudit systems. We propose a family of Bell inequalities that inquire correlations related to the Wigner negativity of stabilizer states under the adjoint action of a generalization of the qubit $\pi/8$ gate, which, in the bipartite case, is an abstraction of the CHSH inequality. The classical bound is simple to compute, and a specified stabilizer state maximally violates the inequality among all qudit states based on the Wigner negativity and an inequality between the 1-norm and the maximum norm. The Bell operator not only serves as a measure for the singlet fraction but also quantifies the volume of Wigner negativity. Furthermore, we give deterministic Bell violations, as well as violations with a constant number of measurements, for the Bell state relying on operators innate to higher-dimensional systems than the qudit at hand.
@@
As a canonical generalization of the classical bit as a unit of information, quantum information theory extensively studied two-level systems, now called qubits. A generalization to higher dimensions, systems of $d$ levels are called qudits. Although qubits are typically easier to describe, qudits can be more efficient for some quantum information tasks: apart from the theoretical interest in deriving no-go theorems in quantum information shaari2008blind , PhysRevA.78.012353 , PhysRevLett.102.120501 , qudit-based algorithms for Byzantine agreement Fitzi2001 and secret sharing in graph states Markham2008 , Kashefi_2009 , Keet2010 appear to surpass schemes known for qubits Cleve1999 . Despite any advantage over qubit systems, quantum information with qudits broadens the range of physical platforms where quantum information tasks can be implemented as many quantum systems are naturally higher dimensional.
We study nonlocality in systems of two or more qudits, referring to quantum correlations that cannot be explained by any local hidden variable model bell1964einstein , brunner2014bell . Previous studies have been conducted to expose nonlocality in high-dimensional systems with Bell states Collins2001 , Liang2009 , Ji2008 , salavrakos2017bell , Greenberger-Horne-Zeilinger (GHZ) states Cerf2001 , and graph states Tang2013 . However, the studied Bell inequalities are hard to analyze, and it is often unclear if they expose genuinely high-dimensional properties. Instead, they resort to properties that are inherent to systems of lower, higher or composite dimension by splicing qudits into subsystems, e.g. subsets of levels, and exploiting effects inherent to the subsystems. In three-dimensional systems, studies have been conducted in Lawrence2017 , Kaszlikowski2002 , Kaszlikowski2002-2 , Acin2004 , Gruca2011 , Li2011 , Mackeprang2023 . Notably, Kaniewski et al. Kaniewski2018 investigate Bell nonlocality based on intrinsic properties of $d$-level systems and even introduce self-testing mutually unbiased bases for the qutrit case. In Howard2015bell , the author uses a known Bell operator Ji2008 to find non-locality with magic states using tools from Howard2012 , which we also employ in this work.
@@
For any odd prime dimension $d$, we construct a family of Bell operators on two qudits in Eq. (10) that relates to Gross’s Wigner function and provides a measure for the singlet fraction as well as for the volume of Wigner negativity. The Bell operators consist of qudit Pauli operators rotated by a generalization of the qubit $T$- or $\pi/8$-gate introduced in Howard2012 , which we call unitary cube operators. Mapping the Pauli group to the qudit Clifford group appleby2005symmetric , Appleby2009 , unitary cube operators cause Wigner negativity in stabilizer states when these states are rotated by them. A corresponding Bell inequality is violated by the Bell state and is a natural extension of the Clauser-Horne-Shimony-Holt (CHSH) inequality for qubits PhysRevLett.23.880 , which is similarly a linear combination of Pauli operators under the local rotation of the $T$-gate. We can then reduce the number of measurement settings constituting the Bell operator by only testing the stabilizer elements of the Bell state with the same implications as before. While the underlying mechanisms for Bell violations are equivalent to the ones Howard2015bell , our main contribution is a broader generalization of the qubit CHSH inequality adaptable to operators beyond the unitary cube operators, as well as a general insight into the origins of non-locality in qudit systems. Lastly, this Bell operator can be adapted to any entangled qudit stabilizer state that maximally violates a corresponding Bell inequality, though it does not detect genuine multipartite entanglement. For $d>3$, unitary cube operators are inherently defined in the qudit system, that is, their eigenvalues are $d^{\text{th}}$ roots of unity, called characters of the finite field wiki:Dirichlet_character . Alternatively, we show strong nonlocal violations for Bell states in $d$ dimension with operators whose spectrum contains characters of yet higher-dimensional systems.
@@
The prime setting for nonlocality, in a Bell scenario for $n$ parties, each party performs a local measurement $O_{m_{i}}$ depending on the local measurement setting $m_{i}$. Physical models can pose restrictions on the general probability $p({\bm{a}}|{\bm{m}})=p((a_{i})_{i}|(m_{i})_{i})$ of the outcome ${\bm{a}}$ for the measurement settings ${\bm{m}}$. In the quantum model, the probability distribution is derived from Born’s rule $p({\bm{a}}|{\bm{m}})=\mathrm{Tr}\left(\Pi^{{\bm{m}}}_{a^{\ast}}\rho\right)$ for a $n$-partite quantum state $\rho$ and projectors $\Pi^{{\bm{m}}}_{a^{\ast}}$ of the joint measurement $\bigotimes_{i=1}^{n}O_{m_{i}}$ with outcome $a^{\ast}=\prod_{i=1}^{n}a_{i}$. The (classical) local hidden variable (lhv) model is more restrictive with probability distributions of the form $p({\bm{a}}|{\bm{m}})=\sum_{\lambda}\mu(\lambda)\prod_{i=1}^{n}p_{i}(a_{i}|m_{i},\lambda)\,,$ for probability distributions $\mu$ and $p_{i}$. The separation can be proven with a Bell inequality for a Bell operator $\mathcal{B}=\sum_{{\bm{m}}}c_{{\bm{m}}}\bigotimes_{i=1}^{n}O_{m_{i}}\,,$ with coefficients $c_{{\bm{m}}}$. If an lhv model describes the probability distribution $p$, then
@@
The qudit Pauli group $\mathcal{P}$ can be decomposed into $d+1$ subgroups of commuting operators $\mathcal{G}_{r}=\{T_{(s,sr)}\,,s\in\mathbb{Z}_{d}\}\subset\mathcal{P}$ for $r=0,\dots,d-1$ and $\mathcal{G}_{d}=\{T_{(0,z)}\,,z\in\mathbb{Z}_{d}\}$. The eigenbases of two elements of distinct groups are mutually unbiased, i.e., their respective eigenvector’s inner product is of magnitude $1/\sqrt{d}$. Taking an element from each commutative group defines a complete set of mutually unbiased bases through their eigenvectors.
@@
generate Gross’s Wigner function
$W_{u}(\rho):=d^{-1}\mathrm{Tr}\left(A_{u}\rho\right)\,,$ (4)
a quasi-probability distribution that is real-valued and normalized, but can have negative values Gross2006 .
For a system of $n$ qudits, we denote $\bm{u}=({\bm{u}_{x}},{\bm{u}_{z}})\in\mathbb{Z}_{d}^{2n}$, and the operator $T_{\bm{u}}=\bigotimes_{i=1}^{n}T_{((u_{i})_{x},(u_{i})_{z})}$, which holds similarly for $A_{\bm{u}}$ as well as for the Wigner function $W_{\bm{u}}$for multiple qudits. In this context, we also use the inner product ${\bm{u}}{\bm{v}}=\sum_{i=1}^{n}u_{i}v_{i}$, which also extends to the symplectic form $[\bm{u},\bm{v}]=\bm{u}_{z}\bm{v}_{x}-\bm{u}_{x}\bm{v}_{z}$.
@@
For $d>3$, it is $f(k)=\nu_{k}=12^{-1}k(\gamma+k(6z+(2k+3)\gamma))+\epsilon k$ with $z,\epsilon\in\mathbb{Z}_{d}$ and $\gamma\in\mathbb{Z}_{d}^{\ast}=\mathbb{Z}_{d}\setminus\{0\}$. We discuss the case $d=3$ later together with non-character unitary operators. The coefficients are such that $U^{\phantom{\dagger}}_{\nu}XU^{\dagger}_{\nu}=\omega^{\epsilon}XZ^{z}\sum_{k\in\mathbb{Z}_{d}}\omega^{\gamma/2}|k\rangle\langle k|$.
@@
For completeness, note that $a_{1}(u_{1},u_{2})=\epsilon+(u_{1})_{z}+(u_{2})_{z}+z(u_{1})_{x}+2^{-1}\gamma((u_{1})_{x}^{2}-(u_{1})_{x}+6^{-1})$ and $a_{3}=24^{-1}\gamma$, such that, independently of $u_{1}$ and $u_{2}$, $a_{1}$ can take any value in $\mathbb{Z}_{d}$ and $a_{3}$ any value in $\mathbb{Z}_{d}^{\ast}$. For $(u_{1})_{x}=(u_{2})_{x}$, the value of $W^{\nu}_{u_{1},u_{2}}:=W_{u_{1},u_{2}}(|\Phi_{\nu}\rangle\langle\Phi_{\nu}|)$ is determined by the sum of characters with a third-order odd polynomial as an argument. Although such polynomials are difficult to analyze for general odd prime $d$, a simple brute-force search finds values $a_{1},a_{3}$ such that $W^{\nu}_{u_{1},u_{2}}<0$ for any $u_{1}$ and $u_{2}$. The negativity of $W^{\nu}_{u_{1},u_{2}}$ is pivotal in achieving a Bell violation with the Bell operator in Eq. (10).
@@
Table 1: Extremal values of the Wigner function $W^{\nu}_{v_{1},v_{2}}$ in Eq. (6) and Eq. (17), maximal values of its negativity volume $N[\Phi_{\nu}]$, all found through an exhaustive search, as well as values of $C_{\min}$ (Eq. (8)) for small prime dimension $d$. For $d\leq 7$, we saturate the bound of Eq. (9) and achieve $C_{\min}=d^{3}\min W^{\nu}_{v_{1},v_{2}}$. * For $d=3$, we instead use $\chi(f)=1+2\cos\left(8\pi/9\right)$ as we allow for cube roots of the character. ** For $d=2$, we use the Wigner function adapted to qubits for the Bell state and the $T$-gate in Eq. (24). Its minimum value is $-1/\sqrt{2}\approx-0.707$, corresponding to the Tsirelson bound Cirelson1980 . Anyway, the Bell operator (10) and analysis thereof is not apt to qubits.
@@
since $g_{0}=0$ and $a=(d\pm 1)/2$ minimizes $\Re(\omega^{a})$. For comparison, numerical values of $\min_{g}\chi(g)$ for small odd prime numbers $d$ are in Tab. 1 which differ from $W_{\min}=\min_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{v_{1},v_{2}}$ for all $d>7$. For $d\rightarrow\infty$, $C_{\mathrm{min}}\rightarrow 2-d$ while $W_{\min}\geq-2/d^{2}\sqrt{d}$ from Eq. (7). To estimate if $C_{\mathrm{min}}$ is achievable by a polynomial $g$, we study the value set $g(\mathbb{Z}_{d})$ of a polynomial over finite fields. Specifically, $C_{\mathrm{min}}$ requires $g$ to have a value set of three elements. The value set of any polynomial has the lower bound
@@
To construct a Bell inequality for qudit stabilizer states, we measure the operators that make up the Wigner function $W^{\nu}_{v_{1},v_{2}}$ to exploit the Wigner negativity of the stabilizer states under unitary cube operators. To highlight the negative values, a favorable coefficient distinguishes the negative values from the positive ones. The head-on candidate is the Wigner function itself, resulting in the Bell operator
$\mathcal{B}_{\nu}=\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{v_{1},v_{2}}\phantom{\cdot}U^{\phantom{\dagger}}_{\nu}A_{v_{1}}U^{\dagger}_{\nu}\otimes A_{v_{2}}\,.$ (10)
@@
While we write the Bell operator using the operators $A_{v}$, the measurements are defined in the measurements in terms of the Pauli operators with projectors $\Pi^{(v)}_{m}=d^{-1}\sum_{k\in\mathbb{Z}_{d}}\omega^{mk}T_{v}^{k}$ of outcome $m\in\mathbb{Z}_{d}$ and $v\in\mathbb{Z}_{d}^{2}$. Although the operator comprises of $d^{4}$ local operators $A_{v}$, only $(d+1)^{2}$ measurements corresponding to mutually unbiased bases for each qudit are necessary. The Bell operator’s expectation value for the Bell state is
@@
$\mathrm{Tr}\left(\mathcal{B}_{\nu}|\Phi\rangle\langle\Phi|\right)=\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}d^{2}(W^{\nu}_{v_{1},v_{2}})^{2}=1\,,$ (11)
since the rotated Bell state is a pure state. The expectation value of the Bell operator for any bipartite state $\sigma$ is a measure for the singlet fraction

### v3

# Bell Nonlocality from Wigner Negativity in Qudit Stabilizer States

@@ Abstract
Nonlocality is an essential concept that distinguishes quantum from classical models and has been extensively studied in systems of qubits. For higher-dimensional systems, certain results for their two-level counterpart, like Bell violations with stabilizer states and Clifford operators, do not generalize. On the other hand, similar to continuous variable systems, Wigner negativity is necessary for nonlocality in qudit systems. We show how a bipartite entangled qudit state can serve as a witness for nonlocality when it exhibits Wigner negativity. Additionally, we propose a new generalization of the CHSH inequality for qudits by inquiring correlations related to the Wigner negativity of stabilizer states under the adjoint action of a generalization of the qubit $\pi/8$-gate. A specified stabilizer state maximally violates the inequality among all qudit states based on its Wigner negativity. The Bell operator not only serves as a measure for the singlet fraction but also quantifies the volume of Wigner negativity. Furthermore, we give deterministic Bell violations, as well as violations with a constant number of measurements, for the Bell state relying on operators innate to higher-dimensional systems than the qudit at hand.
@@
As a canonical extension of the classical bit as a unit of information, quantum information theory extensively studied two-level systems, now called qubits. A generalization to higher dimensions, systems of $d$ levels, are called qudits. Although qubits are typically easier to describe, qudits can be more efficient for some quantum information tasks: apart from the theoretical interest in deriving no-go theorems in quantum information shaari2008blind , PhysRevA.78.012353 , PhysRevLett.102.120501 , qudit-based algorithms for Byzantine agreement Fitzi2001 and secret sharing in graph states Markham2008 , Kashefi_2009 , Keet2010 appear to surpass schemes known for qubits Cleve1999 . Despite any advantage over qubit systems, quantum information with qudits broadens the range of physical platforms where quantum information tasks can be implemented as many quantum systems are naturally higher dimensional.
We study nonlocality in systems of two or more qudits, referring to quantum correlations that cannot be explained by any local-hidden-variable model bell1964einstein , brunner2014bell . Previous studies have been conducted to expose nonlocality in high-dimensional systems with Bell states Collins2001 , Liang2009 , Ji2008 , salavrakos2017bell , Greenberger-Horne-Zeilinger (GHZ) states Cerf2001 , and graph states Tang2013 . However, the studied Bell inequalities are hard to analyze, and it is often unclear if they expose genuinely high-dimensional properties. Instead, they resort to properties that are inherent to systems of lower, higher or composite dimension by regarding subsystems, e.g. subsets of levels, and exploiting effects inherent to the subsystems. In three-dimensional systems, studies have been conducted in Lawrence2017 , Kaszlikowski2002 , Kaszlikowski2002-2 , Acin2004 , Gruca2011 , Li2011 , Mackeprang2023 . Notably, Kaniewski et al. Kaniewski2018 investigate Bell nonlocality based on intrinsic properties of $d$-level systems and even introduce self-testing mutually unbiased bases for the qutrit case. In Howard2015bell , the author uses a known Bell operator Ji2008 to find nonlocality with magic states using tools from Howard2012 , which we also employ in this work.
@@
For any odd prime dimension $d$, we first show in Eq. (11) how certain bipartite entangled states that exhibit Wigner negativity can be witnesses for nonlocality. Then, we construct a family of Bell operators on two qudits in Eq. (16) that relates to Gross’s Wigner function and provides a measure for the singlet fraction as well as for the volume of Wigner negativity. The Bell operators consist of qudit Pauli operators rotated by a generalization of the qubit $T$- or $\pi/8$-gate introduced in Howard2012 , which we call unitary cube operators. Mapping the Pauli group to the qudit Clifford group appleby2005symmetric , Appleby2009 , unitary cube operators cause Wigner negativity in stabilizer states when these states are rotated by them. A corresponding Bell inequality is violated by the Bell state and is a natural extension of the Clauser-Horne-Shimony-Holt (CHSH) inequality for qubits PhysRevLett.23.880 , which is similarly a linear combination of Pauli operators under the local rotation of the $T$-gate. We can then reduce the number of measurement settings constituting the Bell operator by only testing the stabilizer elements of the Bell state with the same implications as before. While the underlying mechanisms for Bell violations are equivalent to the ones Howard2015bell , our main contribution is a broader generalization of the qubit CHSH inequality for qudits adaptable to operators beyond the unitary cube operators, as well as a general insight into the origins of nonlocality in qudit systems. Lastly, this Bell operator can be adapted to any entangled qudit stabilizer state that maximally violates a corresponding Bell inequality, though it does not detect genuine multipartite entanglement beyond three parties. For $d>3$, unitary cube operators are inherently defined in the qudit system, that is, their eigenvalues are $d^{\text{th}}$ roots of unity, called characters of the finite field wiki:Dirichlet_character . Alternatively, we show strong nonlocal violations for Bell states in $d$ dimension with operators whose spectrum contains characters of yet higher-dimensional systems.
@@
The prime setting for nonlocality, in a Bell scenario for $n$ parties, each party performs a local measurement $O_{m_{i}}$ depending on the local measurement setting $m_{i}$. Physical models can pose restrictions on the general probability $p({\bm{a}}|{\bm{m}})=p((a_{i})_{i}|(m_{i})_{i})$ of the outcome ${\bm{a}}$ for the measurement settings ${\bm{m}}$. In the quantum model, the probability distribution is derived from Born’s rule $p({\bm{a}}|{\bm{m}})=\mathrm{Tr}\left(\Pi^{{\bm{m}}}_{a^{\ast}}\rho\right)$ for a $n$-partite quantum state $\rho$ and projectors $\Pi^{{\bm{m}}}_{a^{\ast}}$ of the joint measurement $\bigotimes_{i=1}^{n}O_{m_{i}}$ with outcome $a^{\ast}=\prod_{i=1}^{n}a_{i}$. The (classical) local-hidden-variable (lhv) model is more restrictive with probability distributions of the form $p({\bm{a}}|{\bm{m}})=\sum_{\lambda}\mu(\lambda)\prod_{i=1}^{n}p_{i}(a_{i}|m_{i},\lambda)\,,$ for probability distributions $\mu$ and $p_{i}$. The separation can be proven with a Bell inequality for a Bell operator $\mathcal{B}=\sum_{{\bm{m}}}c_{{\bm{m}}}\bigotimes_{i=1}^{n}O_{m_{i}}\,,$ with coefficients $c_{{\bm{m}}}$. If an lhv model describes the probability distribution $p$, then
@@
The qudit Pauli group $\mathcal{P}$ can be decomposed into $d+1$ subgroups of commuting operators $\mathcal{G}_{r}=\{T_{(s,sr)}\,,s\in\mathbb{Z}_{d}\}\subset\mathcal{P}$ for $r=0,\dots,d-1$ and $\mathcal{G}_{d}=\{T_{(0,z)}\,,z\in\mathbb{Z}_{d}\}$. The eigenbases of two elements of distinct groups are mutually unbiased, i.e., their respective eigenvector’s inner product is of magnitude $1/\sqrt{d}$. Taking an element from each commutative group defines a complete set of mutually unbiased bases through their eigenvectors. Any qudit state has a unique representation $\rho=\frac{1}{d}\sum_{u\in\mathbb{Z}_{d}^{2}}\chi_{u}(\rho)T_{u}$ via its complex characteristic function
$\chi_{u}(\rho)=\mathrm{Tr}(\rho T_{-u})\,.$ (3)
@@
generate Gross’s Wigner function, $W_{u}(\rho):=d^{-1}\mathrm{Tr}\left(A_{u}\rho\right)$, a quasi-probability distribution that is real-valued and normalized, but can have negative values Gross2006 . Any qudit state can be written as $\rho=\sum_{u\in\mathbb{Z}_{d}^{2}}W_{u}(\rho)A_{v}$, and
$W_{u}(\rho)=\frac{1}{d}\sum_{v\in\mathbb{Z}_{d}^{2}}\omega^{-[u,v]}\chi_{v}(\rho)\,,$ (5)
For a system of $n$ qudits, we denote $\bm{u}=({\bm{u}_{x}},{\bm{u}_{z}})\in\mathbb{Z}_{d}^{2n}$, and the operator $T_{\bm{u}}=\bigotimes_{i=1}^{n}T_{((u_{i})_{x},(u_{i})_{z})}$, similarly for $A_{\bm{u}}$, which also define analogously the Wigner function $W_{\bm{u}}(\sigma)$ and characteristic function $\chi_{\bm{u}}(\sigma)$ for a state $\sigma$ on multiple qudits. In this context, we also use the inner product ${\bm{u}}\cdot{\bm{v}}=\sum_{i=1}^{n}u_{i}v_{i}$, which also extends to the symplectic form $[\bm{u},\bm{v}]=\bm{u}_{z}\bm{v}_{x}-\bm{u}_{x}\bm{v}_{z}$.
@@
For $d>3$, it is $f(k)=\nu_{k}=12^{-1}k(\gamma+k(6z+(2k+3)\gamma))+\epsilon k$ with $z,\epsilon\in\mathbb{Z}_{d}$ and $\gamma\in\mathbb{Z}_{d}^{\ast}=\mathbb{Z}_{d}\setminus\{0\}$. We discuss the case $d=3$ later together with non-character unitary operators. The coefficients are such that $U^{\phantom{\dagger}}_{\nu}XU^{\dagger}_{\nu}=\omega^{\epsilon}XZ^{z}\sum_{k\in\mathbb{Z}_{d}}\omega^{2^{-1}\gamma k^{2}}|k\rangle\langle k|$.
@@
Note that $a_{1}(u_{1},u_{2})=\epsilon+(u_{1})_{z}+(u_{2})_{z}+z(u_{1})_{x}+2^{-1}\gamma((u_{1})_{x}^{2}-(u_{1})_{x}+6^{-1})$ and $a_{3}=24^{-1}\gamma$, such that, independently of $u_{1}$ and $u_{2}$, $a_{1}$ can take any value in $\mathbb{Z}_{d}$ and $a_{3}$ any value in $\mathbb{Z}_{d}^{\ast}$. For $(u_{1})_{x}=(u_{2})_{x}$, the value of $W^{\nu}_{u_{1},u_{2}}:=W_{u_{1},u_{2}}(|\Phi_{\nu}\rangle\langle\Phi_{\nu}|)$ is determined by the sum of characters with a third-order odd polynomial as an argument. Although such polynomials are difficult to analyze for general odd prime $d$, a simple brute-force search finds values $a_{1},a_{3}$ such that $W^{\nu}_{u_{1},u_{2}}<0$ for any $u_{1}$ and $u_{2}$. The negativity of $W^{\nu}_{u_{1},u_{2}}$ is pivotal in achieving a Bell violation with the Bell operator in Eq. (11).
@@
Table 1: Extremal values of the Wigner function $W^{\nu}_{v_{1},v_{2}}$ in Eq. (7) and Eq. (21), maximal values of its negativity volume $N^{\nu}:=N[\Phi_{\nu}]$ and the nonlocal bound $B^{\max}_{\mathrm{lhv}}$ from Eq. (14), all found through an exhaustive search, as well as values of $C_{\min}$ (Eq. (9)) for small prime dimension $d$. For $d\leq 7$, we saturate the bound of Eq. (10) and achieve $C_{\min}=d^{3}\min W^{\nu}_{v_{1},v_{2}}$. * For $d=3$, we instead use $\chi(f)=1+2\cos\left(8\pi/9\right)$ as we allow for cube roots of the character. ** For $d=2$, we use the Wigner function adapted to qubits for the Bell state and the $T$-gate in Eq. (28). Its minimum value is $-1/\sqrt{2}\approx-0.707$, corresponding to Tsirelson’s bound Cirelson1980 . Anyway, the Bell operator (11) and analysis thereof is not apt to qubits.
@@
since $g_{0}=0$ and $a=(d\pm 1)/2$ minimizes $\Re(\omega^{a})$. For comparison, numerical values of $\min_{g}\chi(g)$ for small odd prime numbers $d$ are in Tab. 1 which differ from $W_{\min}=\min_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{u_{1},u_{2}}$ for all $d>7$. For $d\rightarrow\infty$, $C_{\mathrm{min}}\rightarrow 2-d$ while $W_{\min}\geq-2/d^{2}\sqrt{d}$ from Eq. (8). To estimate if $C_{\mathrm{min}}$ is achievable by a polynomial $g$, we study the value set $g(\mathbb{Z}_{d})$ of a polynomial over finite fields. Specifically, $C_{\mathrm{min}}$ requires $g$ to have a value set of three elements. The value set of any polynomial has the lower bound
@@
We propose a novel Bell operator in which a specific state $\sigma$ becomes first a witness for contextuality and Wigner negativity and ,then, for nonlocality. The operator, corresponding to a entangled two-qudit state $\sigma$ is
$\mathcal{B}[\sigma]=\sigma=\frac{1}{d^{2}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}(\sigma)\,T_{u_{1}}\otimes T_{u_{2}}\,,$ (11)
with the state’s characteristic function $\chi_{u_{1},u_{2}}(\sigma)=\mathrm{Tr}(\sigma\,T_{-u_{1}}\otimes T_{-u_{2}})$.
@@
The corresponding measurements are defined in terms of projectors $\Pi^{(v)}_{m}=d^{-1}\sum_{k\in\mathbb{Z}_{d}}\omega^{mk}T_{v}^{k}$ of outcome $m\in\mathbb{Z}_{d}$ and $v\in\mathbb{Z}_{d}^{2}$ on the eigenbasis of the Pauli operators. Although the operator comprises of $d^{4}$ local operators $T_{v}$, only $(d+1)^{2}$ measurements corresponding to mutually unbiased bases for each qudit are necessary.
On at least two qudits, a non-contextual (nc) model describes the outcomes of Pauli measurements $T_{u_{1}}\otimes T_{u_{2}}$ that can be performed simultaneously as multiplicative characters of the form $\omega^{[a_{1},u_{1}]+[a_{2},u_{2}]}$ for $a_{1},a_{2}\in\mathbb{Z}_{d}^{2}$ Delfosse2017 . Such a deterministic non-contextual value assignment (nc) can only achieve
@@
$\langle\mathcal{B}[\sigma]\rangle_{\mathrm{nc}^{\ast}}$ $=\frac{1}{d^{2}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}(\sigma)\,\omega^{[a_{1},{u_{1}}]}\omega^{[a_{2},{u_{2}}]}$
$=d^{2}\,W_{a_{1},a_{2}}(\sigma)\,.$ (12)
As a result, any nc model, a convex mixture of deterministic classical value assignment (nc), can maximally achieve
$\left\langle\mathcal{B}[\sigma]\right\rangle_{\mathrm{nc}}\leq d^{2}\max_{a_{1},a_{2}\in\mathbb{Z}_{d}^{2}}|W_{a_{1},a_{2}}(\sigma)|\equiv B_{\mathrm{nc}}^{\max}\,.$ (13)
On the other hand, a pure state state $\sigma$ maximally violates the non-contextual inequality, which it defines via its characteristic function, among all qudit states if $\mathrm{Tr}(\sigma^{2})=1>B_{\mathrm{nc}}^{\max}$. We find with Eq. (31) that non-trivial negativity volume, $N[\Psi]>0$, is necessary for such a violation. As a result, Eq. (11) defines a family of inequalities based on Wigner negative states that are witnesses of contextuality if $1>d^{2}\max_{a_{1},a_{2}\in\mathbb{Z}_{d}^{2}}|W_{a_{1},a_{2}}[\sigma]|$. This specifically exploit the equivalence of Wigner negativity and contextuality with Pauli measurements Delfosse2017 .
If we spatially separate the qudits, and the Pauli measurements cannot be performed simultaneously, a local hidden variable (lhv) model is less restrictive than the non-contextual one, and instead can assign a deterministic value $\omega^{\alpha_{r}}$ with $\alpha_{r}\in\mathbb{Z}_{d}$, $r=0,\dots,d$, for each of the $d+1$ measurement settings, independently. Then, any lhv model can maximally attain
$\langle\mathcal{B}[\sigma]\rangle_{\mathrm{lhv}}\leq\max_{\alpha,\beta\in\mathbb{Z}_{d}^{d+1}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}[\sigma]\,\omega^{\alpha_{r(u_{1})}+\beta_{r(u_{2})}}\equiv B^{\max}_{\mathrm{lhv}}\,.$ (14)
In terms of the Wigner function, the Bell bound is
$B^{\max}_{\mathrm{lhv}}=\max_{\alpha,\beta\in\mathbb{Z}_{d}^{d+1}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W_{u_{1},u_{2}}[\sigma]F(\alpha)_{u_{1}}F(\beta)_{u_{1}}\,.$ (15)
with the discrete Fourier transform of $\omega^{a_{r(u)}}$, $F(a)_{u}=\sum_{v\in\mathbb{Z}_{d}^{2}}\omega^{[u,v]+\alpha_{r(v)}}/d$. The values of $F(\alpha)_{u}$ range between $-1,0,\dots,d$, the sum of entries is $\sum_{u\in\mathbb{Z}_{d}^{2}}F(\alpha)_{u}=d$, and the vector’s norm is equally $\|F(\alpha)\|_{2}=d$. For an $\alpha\in\mathbb{Z}_{d}^{d+1}$ such that $F(\alpha)_{v}=d$ for some $v\in\mathbb{Z}_{d}^{2}$, it is $\alpha_{r(u)}=[a,u]$ for some $a\in\mathbb{Z}_{d}^{2}$, and $F(\alpha)_{w}=0$ for all $w\in\mathbb{Z}_{d}^{2}\setminus\{v\}$, which is precisely, the non-contextual value assignment used for the Wigner negativity witness or non-contextual inequality (13). To evaluate the Bell bound (15) further, the Cauchy-Schwarz inequality gives $B^{\max}_{\mathrm{lhv}}[\sigma]\leq\|W[\sigma]\|_{2}\|F(\alpha)\|_{2}\|F(\beta)\|_{2}\leq d^{2}\|W[\sigma]\|_{2}$. To observe a Bell violation, it is necessary that $B_{\mathrm{lhv}}[\sigma]<1$, which requires at least $\|W[\sigma]\|_{2}\leq 1$. Since the non-contextual value assignments are a special case of the deterministic lhv models’, it is $B^{\max}_{\mathrm{nc}}\leq B^{\max}_{\mathrm{lhv}}$ such that Wigner negativity of the pure resource state $\sigma$ as well as the state used for the quantum realization is again essential for nonlocality. Whether the state defines a notable Bell inequality beyond these criteria requires an analysis of the state’s characteristic or Wigner function. Bell states rotated by local unitary cube operators $|\Phi_{\nu}\rangle$, also employed in Howard2015bell , are prime candidates to define a Bell inequality in terms of their characteristic function and Bell bounds in Tab. 1.
Qudit stabilizer states, however, do not exhibit Wigner negativity. To construct a Bell inequality for qudit stabilizer states, we measure the operators that make up the Wigner function $W^{\nu}_{v_{1},v_{2}}$ to exploit the Wigner negativity of the stabilizer states under unitary cube operators. To highlight the negative values, a favorable coefficient distinguishes the negative values from the positive ones, the head-on candidate being the Wigner function itself. For the qudit Bell state the resulting operator is $\mathcal{B}_{\nu}=\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{v_{1},v_{2}}\phantom{\cdot}U^{\phantom{\dagger}}_{\nu}A_{v_{1}}U^{\dagger}_{\nu}\otimes A_{v_{2}}$. In terms of qudit Pauli operators, it is
$\mathcal{B}_{\nu}=\frac{1}{d^{2}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi^{\nu}_{u_{1},u_{2}}\,U^{\phantom{\dagger}}_{\nu}T_{u_{1}}U^{\dagger}_{\nu}\otimes T_{u_{2}}\,,$ (16)
with the characteristic function $\chi^{\nu}_{u_{1},u_{2}}=\langle\Phi_{\nu}|T_{-u_{1}}\otimes T_{-u_{2}})|\Phi_{\nu}\rangle$. The non-contextual bounds arise for the same reasons as those underlying the inequality (13) such that $\left\langle\mathcal{B}_{\nu}\right\rangle_{\mathrm{nc}}\leq d^{2}\max_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}|W^{\nu}_{v_{1},v_{2}}|\equiv B_{\mathrm{nc}}^{\max}\,.$ The Bell operator’s expectation value for the Bell state is $\mathrm{Tr}\left(\mathcal{B}_{\nu}|\Phi\rangle\langle\Phi|\right)=\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}d^{2}(W^{\nu}_{v_{1},v_{2}})^{2}=1\,,$ since the rotated Bell state is a pure state. Moreover, we show in Eq. (26) that the expectation value of the Bell operator for any bipartite state $\sigma$ is a measure for the singlet fraction